OpenStudy (anonymous):

Can someone please help me with this: A wheel rotates with a constant angular acceleration of 3.45 rad/s2. Assume the angular speed of the wheel is 1.90 rad/s at ti = 0. (a) Find the angle through which the wheel rotates between t = 2.00 s and t = 3.80 s. (b) Find the angular speed when t = 3.80 s. (c) What is the magnitude of the angular speed FIVE revolutions following t = 3.80 s? I need help with Part C.

OpenStudy (anonymous):

\[\theta = \omega _{i}t +\frac{ 1 }{ 2 }\alpha t^2\]

OpenStudy (anonymous):

\[\theta(3.8) -\theta(2) = \omega _{i}(3.8) +\frac{ 1 }{ 2 }\alpha (3.8)^2 -(\omega _{i}(2) +\frac{ 1 }{ 2 }\alpha (2)^2)\]

OpenStudy (anonymous):

\[\omega (3.8) =\omega _{i} +\alpha (3.8)\]